OpenStudy (anonymous):

if you double a recipe that requries 1, 1/2 of flour, how much flour will you need?

OpenStudy (anonymous):

what is 2* 1 1/2?

OpenStudy (australopithecus):

No, what is 2 * 1 + 1/2

OpenStudy (australopithecus):

Hint: 2/2 = 1 and to add fractions you need to have the same denominator

OpenStudy (australopithecus):

Also notice that 2 = 2/1 and to multiply fractions y/x * d/c = dy/cx

OpenStudy (australopithecus):

so First solve the addition of fractions 1 + 1/2 you can't add fractions with different denominators 1/1 = 1 as 1 goes into 1, one times notice though that 2/2 = 1 as 2 goes into 2, two times thus we can write the problem like this 2/2 + 1/2 when adding fractions add the numerators so we have (2 + 1)/2 thus 3/2 so we have two times 3/2 2 * 3/2 Notice that 2/1 = 2, as 1 goes into 2, two times so we can right the problem like so (2/1)*(3/2) = (2)(3)/(1)(2) = 6/2 is the amount of flour we need 2*(1 + 1/2) = 2*(3/2) = 6/2 or 3 cups